Find the inverse of A = ( 1,3,0) ( 0,4,-6) ( -1,5,7) .


1
Expert's answer
2021-06-07T13:07:09-0400
"A=\begin{pmatrix}
 1 & 3 & 0 \\
 0 & 4 & -6 \\
 -1 & 5 & 7\\
\end{pmatrix}"

Augment the matrix with the identity matrix:


"\begin{pmatrix}
 1 & 3 & 0 & & 1 & 0 & 0 \\
 0 & 4 & -6 & & 0 & 1 & 0 \\
 -1 & 5 & 7 & & 0 & 0 & 1\\
\end{pmatrix}"

"R_3=R_3+R_1"


"\begin{pmatrix}
 1 & 3 & 0 & & 1 & 0 & 0 \\
 0 & 4 & -6 & & 0 & 1 & 0 \\
 0 & 8 & 7 & & 1 & 0 & 1\\
\end{pmatrix}"

"R_2=R_2\/4"


"\begin{pmatrix}
 1 & 3 & 0 & & 1 & 0 & 0 \\
 0 & 1 & -3/2 & & 0 & 1/4 & 0 \\
 0 & 8 & 7 & & 1 & 0 & 1\\
\end{pmatrix}"

"R_1=R_1-3R_2"


"\begin{pmatrix}
 1 & 0 & 9/2 & & 1 & -3/4 & 0 \\
 0 & 1 & -3/2 & & 0 & 1/4 & 0 \\
 0 & 8 & 7 & & 1 & 0 & 1\\
\end{pmatrix}"

"R_3=R_3-8R_2"


"\begin{pmatrix}
 1 & 0 & 9/2 & & 1 & -3/4 & 0 \\
 0 & 1 & -3/2 & & 0 & 1/4 & 0 \\
 0 & 0 & 19 & & 1 & -2 & 1\\
\end{pmatrix}"

"R_3=R_3\/19"


"\begin{pmatrix}
 1 & 0 & 9/2 & & 1 & -3/4 & 0 \\
 0 & 1 & -3/2 & & 0 & 1/4 & 0 \\
 0 & 0 & 1 & & 1/19 & -2/19 & 1/19\\
\end{pmatrix}"

"R_1=R_1-(9\/2)R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & 29/28 & -21/76 & -9/38 \\
 0 & 1 & -3/2 & & 0 & 1/4 & 0 \\
 0 & 0 & 1 & & 1/19 & -2/19 & 1/19\\
\end{pmatrix}"

"R_2=R_2+(3\/2)R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & 29/38 & -21/76 & -9/38 \\
 0 & 1 & 0 & & 3/38 & 7/76 & 3/38 \\
 0 & 0 & 1 & & 1/19 & -2/19 & 1/19\\
\end{pmatrix}"

On the left is the identity matrix. On the right is the inverse matrix.




"\begin{pmatrix}
 1 & 3 & 0 \\
 0 & 4 & -6 \\
 -1 & 5 & 7\\
\end{pmatrix}\begin{pmatrix}
 29/38 & -21/76 & -9/38 \\
 3/38 & 7/76 & 3/38 \\
 1/19 & -2/19 & 1/19\\
\end{pmatrix}"


"=\begin{pmatrix}
 (29+9+0)/38 & (-21+21-0)/76 &(-9+9+0)/38 \\
 (-0+12-12)/38 & (-0+28+48)/76 & (-0+12-12)/38 \\
 (-29+15+14)/38 & (21+35-56)/76 & (9+15+14)/38\\
\end{pmatrix}"


"=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1\\
\end{pmatrix}"

"A^{-1}=\begin{pmatrix}
 29/38 & -21/76 & -9/38 \\
 3/38 & 7/76 & 3/38 \\
 1/19 & -2/19 & 1/19\\
\end{pmatrix}"